OpenStudy (anonymous):

atmospheric pressure. atmospheric pressure P (in kilopascals, kPa) at altitude h (in kilometers, km) is governed by the formula In (P/Po)=-h/k, where K=7 and Po=100 kPa constants. solve equation of P. use part (a) to find the pressure P at an altitude of 4 km

OpenStudy (danjs):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

dont know how to solve it

OpenStudy (danjs):

In (P/Po)=-h/k is given

OpenStudy (anonymous):

yes

OpenStudy (danjs):

they also give you Po and k

OpenStudy (anonymous):

yes

OpenStudy (danjs):

so ln(P/100) = -h / 7

OpenStudy (anonymous):

k=7 and Po=100 kPa

OpenStudy (danjs):

we want to solve this for P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes

OpenStudy (danjs):

recall: ln(a/b) = ln(a) - ln(b)

OpenStudy (anonymous):

yes i remember

OpenStudy (danjs):

ln(P) - ln(100) = -h / 7

OpenStudy (danjs):

ln(p) = -h/7 + ln(100)

OpenStudy (anonymous):

you added In(100) to both sides?

OpenStudy (danjs):

yes recall: \[e ^{\ln(a)}=a\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats the next step

OpenStudy (danjs):

ok

OpenStudy (danjs):

\[e ^{\ln(P)} = e ^{(-h/7 + \ln(100))}\]

OpenStudy (anonymous):

oh the e value

OpenStudy (danjs):

recall: \[e ^{a+b} = e ^{a}*e ^{b}\]

OpenStudy (danjs):

\[P = e ^{-h/7} * e ^{\ln(100)}\]

OpenStudy (danjs):

\[P = 100*e ^{-h/7} \]

OpenStudy (danjs):

For h = 4 km, What is P?

OpenStudy (danjs):

you get everything?

OpenStudy (anonymous):

oh sorry no

OpenStudy (anonymous):

Po=100 kPa

OpenStudy (danjs):

Yes, we want to find P when h = 4 km

OpenStudy (danjs):

we solved for P in terms of h above

OpenStudy (anonymous):

oh for part A the book has e^-kh

OpenStudy (danjs):

\[\ln(\frac{ P }{ 100 })=-h/k\] solving for P, \[P = 100 * e ^{\frac{ -h }{ 7 }}\] now you are given h = 4

OpenStudy (anonymous):

i need help with another one

OpenStudy (danjs):

\[P = 100 * e ^{\frac{ -4 }{ 7 }}\] \[P \approx 56.47 kPa\]

OpenStudy (anonymous):

i got that answer

OpenStudy (danjs):

k

OpenStudy (anonymous):

i have a question

OpenStudy (danjs):

can you open a new thread and tag me please

OpenStudy (anonymous):

where u do that

OpenStudy (danjs):

@klogan

OpenStudy (danjs):

yeh it is just type @DanJS